OpenStudy (anonymous):

1.What percent of 65 is 13 ? 2.25% of what is 13 ? 3.What is 35% of 180 ? 4.What percent of 34 is 17 ? ----------------------- a.50 b.20 c.52 d.63 Which One?

OpenStudy (anonymous):

Please Help D:

OpenStudy (anonymous):

So for: a)(13/65)*100= b) 13/(.25) = c) .35*180= d) (17/34)*100=